OpenStudy (anonymous):

1. 3t+5=2 1. 3t+5=2 @Mathematics

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

3t+5=2 cancel out the 5 3t=2-5=-3 cancel the 3 t=-3/3=-1 t=-1